Give the equation of a conic which is symmetric to the line x+2 = 0.
1
Expert's answer
2020-02-19T09:50:52-0500

The general equation of conic section is

"Ax^2 +Bxy +Cy^2+Dx+Ey +F =0\nwhere, A,B,C,D,E AND F are constant"

we know that as the value of constant change then the shape of conic section will also change

we have given line equation as , x+2=0 about this line we have to find the equation of conic section

take, x=-2 now when we compare it with

"A(x+2)^2 +B(x+2)y +Cy^2+D(x+2) +Ey +F =0"


"A(x^2 +2x+4)+ Bxy +2By +Cy^2 +2Dx+2D +Ey+F=0"

"Ax^2 +Bxy+ Cy^2+ (2A+2D)x + (2B+E)y + (4A+2D+F)=0"

the above equation is symmetrical about the line x+2=0